# Constraining the Cosmic-ray Energy Based on Observations of Nearby Galaxy Clusters by LHAASO

### v1

@@ 4 Discussion and Conclusion
Based on the aforementioned constraints on $\gamma$-rays, we can place limits on the energy budget of CRp in galaxy clusters. LHAASO measured the $\gamma$-ray emissions above $1\leavevmode\nobreak\ \rm TeV$, which correspond to the CRp energy of more than $10\leavevmode\nobreak\ \rm TeV$ (Kafexhiu et al., 2014; Kelner et al., 2006). Thus, the $\gamma$-ray upper limits we derived above can be used to derive the model-independent CRp energy budget in this energy range directly, which is listed in Table 2.
The CRp are accelerated by various shocks and turbulent environments within galaxy clusters and then diffuse throughout the cluster. The large scale of galaxy clusters and their magnetic fields (on the order of a few $\mu$G) endows CRp with an escape timescale comparable to cosmological timescales. Furthermore, due to the low density of the ICM, the cooling timescale for CRp exceeds the Hubble time, signifying that CRp in galaxy clusters retains information about their acceleration and propagation. The total CRp energy in the cluster reflects the efficiency of cosmic-ray acceleration. Therefore, if we assume that the extended $\gamma$-ray radiation from galaxy clusters is due to the collision of CRp with the ICM, the upper limits previously derived can constrain the physics related to CRp acceleration and propagation for these three galaxy clusters.
@@
In scenarios where CRs accelerated within the cluster experience full confinement, the resulting CRp population would continuously accumulate over time rather than reaching a stationary state. Under the assumption of energy-independent hadronic interaction cross sections, the resulting $\gamma$-ray emission spectrum would then reflect this time-integrated CRp population. The energy budget integrated over the entire energy spectrum of CRp can be used to derive $X_{CR}$. However, the confinement of CRs is determined by the properties of turbulence in these regions, which can result in an energy-dependent confinement time (Blasi et al., 2007). The confinement can then be effective only below some critical energy $E_{c}$. Below $E_{c}$, the confinement time is larger than the Hubble time, and the $\gamma$-ray spectrum is the same as the injected/accelerated one. For energies above $E_{c}$, the $\gamma$-ray spectrum is softened due to the escape of higher-energy CRs. In LHAASO observations, the most stringent constraints come from the first several energy bins of WCDA of several $\rm TeV$, which correspond to the CRp energy of dozens of TeV. Thus, if $E_{c}>30\leavevmode\nobreak\ \rm TeV$, the constraints derived from LHAASO can be extrapolated to lower energies according to the acceleration spectrum, which is typically a power law. On the other hand, if $E_{c}<30\leavevmode\nobreak\ \rm TeV$, we expect a softening of the $\gamma$-ray spectrum in the WCDA energy band. Thus, the CRp energy budget above $10\leavevmode\nobreak\ \rm TeV$ we derived here cannot be extrapolated to lower energies to constrain $X_{CR}$.
@@
In addition to the uncertain confinement conditions, other significant uncertainties persist, particularly concerning the energy and spatial distributions of CRp and the ICM. Under the assumptions that the CRp energy spectrum follows a power-law distribution and that cosmic rays blow 30 TeV can be completely confined within galaxy clusters, when considering the Diffusive Shock Acceleration (DSA) of CRp, we expect the spatially integrated CRp energy spectrum to be the same as the CRp injection spectrum and to be hard. For example, it could have an index of -2.1 or -2.3. However, a low shock Mach number (Gabici & Blasi, 2003) and propagation effects may lead to a softer spectral index in the local spatial region, such as near the CRp source. Thus, we investigated four scenarios with spectral indices of -2.1, -2.3, -2.5, and -2.7. It is important to note that both theoretical analyses and simulation results generally support a hard spatially integrated CRp energy spectrum (index $>$ -2.5) (Vannoni et al., 2011; Pinzke & Pfrommer, 2010). Therefore, the constraints based on an index of -2.1 or -2.3 are likely more realistic and deserve emphasis.
@@
Here, we assume a spherically symmetric distribution for both CRp and the ICM. Hypothesizing that the CRp density distribution adheres to the same form as the ICM density distribution, this is to maintain consistency with similar previous studies (HESS Collaboration & Aharonian, 2009; Arlen et al., 2012; MAGIC Collaboration et al., 2016; HESS Collaboration et al., 2023; Ackermann et al., 2010; CTA Consortium et al., 2023). If we further assume that all CRp are confined within the galaxy clusters, we can derive $X_{CR}$ and list the values in Tab.2. However, it must be noted that the $X_{CR}$ we derived here is highly model-dependent.
@@
This results are also shown in Figure 2. The blue dashed line in the left panel represents the expected $\gamma$-ray flux. This flux is computed using the parametrization of the $\gamma$-ray production cross-section in $pp$ collisions, as derived in (Kafexhiu et al., 2014). The calculation is carried out under the most stringent constraints, assuming a spectral index of $-2.1$. Under this assumption, the Coma, Perseus, and Virgo clusters constrain the value of $X_{\text{CRp}}$ to be less than approximately $0.029$, $0.012$, and $0.04$, respectively. The total thermal energy within the $R_{500}$ radius of the three galaxy clusters, denoted as $E_{th_{500}}$, can be found in Table A.1. Constraints on $X_{CRp}$ under other assumptions of spectral index are presented in Table 2.
Radio observations of the synchrotron radiation of electrons in galaxy clusters have indirectly constrained the energy density of relativistic protons in the Coma galaxy cluster. Assuming a magnetic field of $2\,\mu$G in the galaxy cluster, the constraints on $X_{\text{CRp}}$ are $X_{\text{CRp}}<0.009\%$, $0.01\%$, and $0.07\%$ based on cosmic-ray index assumptions of 2.1, 2.3, and 2.5, respectively (Reimer et al., 2004). Observations of the magnetic field in Coma cluster indicate that the magnetic field at the center of the it ranges from $3.9$ to $5.4\,\mu$G (Bonafede et al., 2010). This implies that the constraints from radio observations are at least an order of magnitude tighter than $\gamma$-ray constraints, highlighting the complementarity of multi-wavelength approaches. However, our $\gamma$-ray constraints remain complementary, particularly for scenarios with suppressed magnetic fields or non-standard secondary pair cooling (Enßlin, 2002; Brunetti et al., 2013). For comparison, the constraints on $X_{cr}$ from other $\gamma$-ray detectors are presented on the right panel of Figure 2.
@@
Coma pp index=-2.1 0.029 $7.77\times 10^{61}$ $1.96\times 10^{61}$
@@
Perseus pp index=-2.1 0.012 $2.48\times 10^{61}$ $5.98\times 10^{60}$
@@
Virgo pp index=-2.1 0.038 $3.21\times 10^{60}$ $8.33\times 10^{59}$
@@
For further details, please refer to the article Vannoni et al. (2011).
@@
d 95% C.L. upper limit for the total energy of CRp with energy $>$ 10 TeV within the $R_{500}$ radius. The unit is erg.
@@
In addition to the $pp$ interaction of CR protons with ambient gas, other mechanisms in galaxy clusters, such as the IC processes of relativistic electrons or the $p\gamma$ processes of UHECR accelerated by accretion/merging shocks, can also generate $\gamma$-ray radiation. Indeed, galaxy clusters are believed to be the most ideal sites for the acceleration of UHECRs, given their large size and thus long-enough confinement time (Hillas, 1984; Vannoni et al., 2011; Zirakashvili & Ptuskin, 2019; Condorelli et al., 2023). These UHECRs rapidly lose energy through Bethe-Heitler pair production interactions with the Cosmic Microwave Background Radiation (CMBR), generating high-energy electron-positron pairs. Subsequently, these electrons quickly lose energy via the IC and synchrotron processes, leaving a $\gamma$-ray signature near the site of UHECR acceleration. In such cases, the emission would follow the spatial distribution of the shocks and present annulus/ring-type morphologies. However, we found no indications of such $\gamma$-ray emissions in the current LHAASO skymaps.
@@
Based on this physical process, in addition to the previous constraints derived from the $pp$ interaction on CRp at tens of TeV, we can set limits on much higher-energy CRs, reaching up to $\sim 10^{18}-10^{19}$ eV. The $\gamma$-ray emission in this scenario reveals a quite hard spectrum, with a peak in the intrinsic spectrum at $10-100\leavevmode\nobreak\ \rm TeV$. This makes LHAASO an ideal instrument for such a study.
@@
In this work, we compared the upper limit we derived with the calculations in Vannoni et al. (2011), where both linear and nonlinear diffusive shock acceleration (DSA) were considered for a fixed total cosmic-ray energy. By comparing Figures 7 and 13 from this study with the upper limits we provided, we further constrained the UHECRs accelerated in galaxy clusters. The solid and dashed gray lines on the left-hand side of Figure 2 respectively show the expected $\gamma$-ray spectra from linear and nonlinear DSA under the aforementioned physical assumptions, scaled to match our most stringent constraints. It should be noted that the upper limit we derived in this work is the intrinsic $\gamma$-ray spectrum before absorption, and it was also compared with the intrinsic $\gamma$-ray spectra calculated in Vannoni et al. (2011).
@@
For linear DSA, we constrained $X_{CRp}$ to be less than 0.12, 0.08, and 0.12 for the Coma, Perseus, and Virgo clusters, respectively. For nonlinear DSA, these limits were tightened to 0.011, 0.008, and 0.011. Detailed information is also provided in Table 2. The tighter constraints in the non-linear DSA scenario are expected, as shock modification produces a harder CR spectrum (Vannoni et al., 2011). As a result, the available kinetic energy transferred to relativistic particles, is accumulated at the highest energies. These high-energy particles dominate the contribution to the pair production and, ultimately, to the IC $\gamma$-ray emissions of the secondary electrons.
@@
It is noteworthy that the propagation effects of UHECRs cannot be entirely disregarded. While Vannoni et al. (2011) did not consider the propagation and escape processes of UHECRs in galaxy clusters, recent simulation studies (Condorelli et al., 2023) show that approximately 40% of $10^{19}\,\text{eV}$ cosmic rays escape from cluster environments with central magnetic fields of $3\,\mu\text{G}$. This significant escape fraction may substantially modify the predicted $\gamma$-ray spectrum. The non-detection of characteristic annulus/ring-type $\gamma$-ray features in LHAASO observations under the $p\gamma$+IC scenario may imply two possibilities: either the UHECR acceleration efficiency is insufficient to reach the required energy densities or UHECRs ($E\gtrsim 10^{18}\,\text{eV}$) have escaped from the cluster environment. We stress that our current conclusions are conditional on the crucial assumption of effective UHECR confinement within galaxy clusters. A comprehensive consideration of particle escape and its energy-dependent effects will be indispensable for future modeling efforts.
@@
In this field, multi-messenger observations, such as those involving neutrinos, can also provide important information. Moreover, the absorption of neutrinos is negligible, which gives them a distinct advantage at TeV energies and above. For example, Shi et al. (2023) utilized IceCube data to constrain the average baryon loading factor in galaxy cluster populations. However, due to the current sensitivity limitations of neutrino detectors, these constraints are not yet as stringent as those obtained from $\gamma$-ray observations. The forthcoming next-generation neutrino detectors, such as KM3NET(Margiotta, 2014), ICECUBE Gen-2(Aartsen et al., 2021), Trident(Ye et al., 2023) and HUNT(Huang et al., 2023), will undoubtedly provide more informative observations and offer deeper insights into this field.
@@ Appendix A Detailed Information OF THESE GALAXY CLUSTERS
To characterize the thermal gas within the ICM across different galaxy clusters, we employed various spatial models for $n_{e}$ and $P_{e}$, as used in previous studies (HESS Collaboration & Aharonian, 2009; Adam et al., 2021; CTA Consortium et al., 2023; HESS Collaboration et al., 2023). Details are provided in Table A.2. For the Virgo cluster, previous studies focused only on the central region ($<$ 44 kpc). In this study, we adopt the latest fitting results for gas and temperature distributions, including data beyond the $R_{500}$ radius, obtained from Planck observations of the Sunyaev–Zel’dovich (SZ) effect (Ade et al., 2016).
@@
Comaa Density profile: $n_{e}(r)=n_{0}\left[1+\left(\frac{r}{r_{c}}\right)^{2}\right]^{-\frac{3\beta}{2}}$ $n_{0}=3.42\times 10^{-3}\mathrm{\leavevmode\nobreak\ cm}^{-3},r_{\mathrm{c}}=290\ \mathrm{kpc},\beta=0.75$
@@
Virgoc Density profile: $n_{e}(r)=n_{0,1}\left[1+\left(\frac{r}{r_{c,1}}\right)^{2}\right]^{-\frac{3\beta_{1}}{2}}\\ \quad\quad\quad+n_{0,2}\left[1+\left(\frac{r}{r_{c,2}}\right)^{2}\right]^{-\frac{3\beta_{2}}{2}}(r\lesssim 30\ kpc)\\ n_{\mathrm{e}}(r)=\frac{8.5\times 10^{-5}}{(r/\mathrm{Mpc})^{1.2}}\mathrm{\leavevmode\nobreak\ cm}^{-3}(r\gtrsim 250\ kpc)$ $n_{0,1}=8.9\times 10^{-2}\ \mathrm{cm}^{-3},r_{c,1}=4\ \mathrm{kpc},\beta_{1}=1,n_{0,2}=1.9\times 10^{-3}\ \mathrm{cm}^{-3},r_{c,2}=19.4\ \mathrm{kpc},\beta_{2}=0.47$
@@
    For Coma, We adopt the $\beta$ model (Cavaliere & Fusco-Femiano, 1978) for the density profile and a generalized Navarro Frenk White (Nagai et al., 2007) model for pressure radial profile, with parameters taken from Briel et al. (1992) and Planck Collaboration: et al. (2013).
@@
    The functions and parameters used for the Perseus cluster in this study are consistent with those employed in CTA Consortium et al. (2023).
@@
    Ade et al. (2016) combines a double-beta function for the inner region $(r\lesssim 30\ kpc)$ with a power-law function for the outer region $(r\gtrsim 250\ kpc)$ to describe the thermal electron number density distribution in the Virgo cluster. A simple linear interpolation was applied to the intermediate region. We converted the angular units into distance units.

### v2

@@ IV Discussion and Conclusion
Based on the aforementioned constraints on $\gamma$-rays, we can place limits on the energy budget of CRp in galaxy clusters. LHAASO measured the $\gamma$-ray emissions above $1\penalty 10000\ \rm TeV$, which correspond to the CRp energy of more than $10\penalty 10000\ \rm TeV$ (kafexhiuParametrizationGammarayProduction2014a; kelnerEnergySpectraGamma2006d). Thus, the $\gamma$-ray upper limits we derived above can be used to derive the model-independent CRp energy budget in this energy range directly, which is listed in Table 2.
[…]
@@
In scenarios where CRs accelerated within the cluster experience full confinement, the resulting CRp population would continuously accumulate over time rather than reaching a stationary state. Under the assumption of energy-independent hadronic interaction cross sections, the resulting $\gamma$-ray emission spectrum would then reflect this time-integrated CRp population. The energy budget integrated over the entire energy spectrum of CRp can be used to derive $X_{CR}$. However, the confinement of CRs is determined by the properties of turbulence in these regions, which can result in an energy-dependent confinement time (blasiGAMMARAYSCLUSTERS2007). The confinement can then be effective only below some critical energy $E_{c}$. Below $E_{c}$, the confinement time is larger than the Hubble time, and the $\gamma$-ray spectrum is the same as the injected/accelerated one. For energies above $E_{c}$, the $\gamma$-ray spectrum is softened due to the escape of higher-energy CRs. In LHAASO observations, the most stringent constraints come from the first several energy bins of WCDA of several $\rm TeV$, which correspond to the CRp energy of dozens of TeV. Thus, if $E_{c}>30\penalty 10000\ \rm TeV$, the constraints derived from LHAASO can be extrapolated to lower energies according to the acceleration spectrum, which is typically a power law. On the other hand, if $E_{c}<30\penalty 10000\ \rm TeV$, we expect a softening of the $\gamma$-ray spectrum in the WCDA energy band. Thus, the CRp energy budget above $10\penalty 10000\ \rm TeV$ we derived here cannot be extrapolated to lower energies to constrain $X_{CR}$.
@@
In addition to the uncertain confinement conditions, other significant uncertainties persist, particularly concerning the energy and spatial distributions of CRp and the ICM. Under the assumptions that the CRp energy spectrum follows a power-law distribution and that cosmic rays blow 30 TeV can be completely confined within galaxy clusters, when considering the Diffusive Shock Acceleration (DSA) of CRp, we expect the spatially integrated CRp energy spectrum to be the same as the CRp injection spectrum and to be hard. For example, it could have an index of -2.1 or -2.3. However, a low shock Mach number (gabiciNonthermalRadiationClusters2003) and propagation effects may lead to a softer spectral index in the local spatial region, such as near the CRp source. Thus, we investigated four scenarios with spectral indices of -2.1, -2.3, -2.5, and -2.7. It is important to note that both theoretical analyses and simulation results generally support a hard spatially integrated CRp energy spectrum (index >> -2.5) (vannoniAccelerationRadiationUltrahigh2011; pinzkeSimulatingGrayEmission2010). Therefore, the constraints based on an index of -2.1 or -2.3 are likely more realistic and deserve emphasis.
@@
Here, we assume a spherically symmetric distribution for both CRp and the ICM. Hypothesizing that the CRp density distribution adheres to the same form as the ICM density distribution, this is to maintain consistency with similar previous studies (thehesscollaborationConstraintsMultiTeVParticle2009; arlenCONSTRAINTSCOSMICRAYS2012; magiccollaborationDeepObservationNGC2016; collaborationConstrainingCosmicrayPressure2023c; ackermannGeVGAMMARAYFLUX2010; consortiumProspectsGammaRay2023). If we further assume that all CRp are confined within the galaxy clusters, we can derive $X_{CR}$ and list the values in Tab.2. However, it must be noted that the $X_{CR}$ we derived here is highly model-dependent.
@@
This results are also shown in Figure 2. The blue dashed line in the left panel represents the expected $\gamma$-ray flux. This flux is computed using the parametrization of the $\gamma$-ray production cross-section in $pp$ collisions, as derived in (kafexhiuParametrizationGammarayProduction2014a). The calculation is carried out under the most stringent constraints, assuming a spectral index of 2.1-2.1. Under this assumption, the Coma, Perseus, and Virgo clusters constrain the value of $X_{\text{CRp}}$ to be less than approximately $0.029$, $0.012$, and $0.04$, respectively. The total thermal energy within the $R_{500}$ radius of the three galaxy clusters, denoted as $E_{th_{500}}$, can be found in Table A.1. Constraints on $X_{CRp}$ under other assumptions of spectral index are presented in Table 2.
[…]
@@
Coma $p$ index=-2.1 0.029 $7.77\times 10^{61}$ $1.96\times 10^{61}$
@@
Perseus $p$ index=-2.1 0.012 $2.48\times 10^{61}$ $5.98\times 10^{60}$
@@
Virgo $p$ index=-2.1 0.038 $3.21\times 10^{60}$ $8.33\times 10^{59}$
@@
For further details, please refer to the article vannoniAccelerationRadiationUltrahigh2011.
@@
d 95% C.L. upper limit for the total energy of CRp with energy >> 10 TeV within the $R_{500}$ radius. The unit is erg.
@@
In addition to the $pp$ interaction of CR protons with ambient gas, other mechanisms in galaxy clusters, such as the IC processes of relativistic electrons or the $p\gamma$ processes of UHECR accelerated by accretion/merging shocks, can also generate $\gamma$-ray radiation. Indeed, galaxy clusters are believed to be the most ideal sites for the acceleration of UHECRs, given their large size and thus long-enough confinement time (hillasOriginUltraHighEnergyCosmic1984; vannoniAccelerationRadiationUltrahigh2011; zirakashviliCosmicRayAcceleration2019; condorelliImpactGalaxyClusters2023b). These UHECRs rapidly lose energy through Bethe-Heitler pair production interactions with the Cosmic Microwave Background Radiation (CMBR), generating high-energy electron-positron pairs. Subsequently, these electrons quickly lose energy via the IC and synchrotron processes, leaving a $\gamma$-ray signature near the site of UHECR acceleration. In such cases, the emission would follow the spatial distribution of the shocks and present annulus/ring-type morphologies. However, we found no indications of such $\gamma$-ray emissions in the current LHAASO skymaps.
@@
Based on this physical process, in addition to the previous constraints derived from the $pp$ interaction on CRp at tens of TeV, we can set limits on much higher-energy CRs, reaching up to $\sim 10^{18}-10^{19}$ eV. The $\gamma$-ray emission in this scenario reveals a quite hard spectrum, with a peak in the intrinsic spectrum at $10-100\penalty 10000\ \rm TeV$. This makes LHAASO an ideal instrument for such a study.
@@
In this work, we compared the upper limit we derived with the calculations in vannoniAccelerationRadiationUltrahigh2011, where both linear and nonlinear diffusive shock acceleration (DSA) were considered for a fixed total cosmic-ray energy. By comparing Figures 7 and 13 from this study with the upper limits we provided, we further constrained the UHECRs accelerated in galaxy clusters. The solid and dashed gray lines on the left-hand side of Figure 2 respectively show the expected $\gamma$-ray spectra from linear and nonlinear DSA under the aforementioned physical assumptions, scaled to match our most stringent constraints. It should be noted that the upper limit we derived in this work is the intrinsic $\gamma$-ray spectrum before absorption, and it was also compared with the intrinsic $\gamma$-ray spectra calculated in vannoniAccelerationRadiationUltrahigh2011.
@@
For linear DSA, we constrained $X_{CRp}$ to be less than 0.12, 0.08, and 0.12 for the Coma, Perseus, and Virgo clusters, respectively. For nonlinear DSA, these limits were tightened to 0.011, 0.008, and 0.011. Detailed information is also provided in Table 2. The tighter constraints in the non-linear DSA scenario are expected, as shock modification produces a harder CR spectrum (vannoniAccelerationRadiationUltrahigh2011). As a result, the available kinetic energy transferred to relativistic particles, is accumulated at the highest energies. These high-energy particles dominate the contribution to the pair production and, ultimately, to the IC $\gamma$-ray emissions of the secondary electrons.
@@
It is noteworthy that the propagation effects of UHECRs cannot be entirely disregarded. While vannoniAccelerationRadiationUltrahigh2011 did not consider the propagation and escape processes of UHECRs in galaxy clusters, recent simulation studies (condorelliImpactGalaxyClusters2023b) show that approximately 40% of $10^{19}\,\text{eV}$ cosmic rays escape from cluster environments with central magnetic fields of $3\,\mu\text{G}$. This significant escape fraction may substantially modify the predicted $\gamma$-ray spectrum. The non-detection of characteristic annulus/ring-type $\gamma$-ray features in LHAASO observations under the $p\gamma$+IC scenario may imply two possibilities: either the UHECR acceleration efficiency is insufficient to reach the required energy densities or UHECRs ($E\gtrsim 10^{18}\,\text{eV}$) have escaped from the cluster environment. We stress that our current conclusions are conditional on the crucial assumption of effective UHECR confinement within galaxy clusters. A comprehensive consideration of particle escape and its energy-dependent effects will be indispensable for future modeling efforts.
@@
In this field, multi-messenger observations, such as those involving neutrinos, can also provide important information. Moreover, the absorption of neutrinos is negligible, which gives them a distinct advantage at TeV energies and above. For example, shiConstrainingBaryonLoading2023a utilized IceCube data to constrain the average baryon loading factor in galaxy cluster populations. However, due to the current sensitivity limitations of neutrino detectors, these constraints are not yet as stringent as those obtained from $\gamma$-ray observations. The forthcoming next-generation neutrino detectors, such as KM3NET(margiottaKM3NeTDeepseaNeutrino2014), ICECUBE Gen-2(aartsenIceCubeGen2WindowExtreme2021), Trident(yeMulticubickilometreNeutrinoTelescope2023) and HUNT(huangProposalHighEnergy2023), will undoubtedly provide more informative observations and offer deeper insights into this field.
@@ Appendix A Detailed Information OF THESE GALAXY CLUSTERS
To characterize the thermal gas within the ICM across different galaxy clusters, we employed various spatial models for $n_{e}$ and $P_{e}$, as used in previous studies (thehesscollaborationConstraintsMultiTeVParticle2009; adamRayDetectionComa2021; consortiumProspectsGammaRay2023; collaborationConstrainingCosmicrayPressure2023c). Details are provided in Table A.2. For the Virgo cluster, previous studies focused only on the central region (<< 44 kpc). In this study, we adopt the latest fitting results for gas and temperature distributions, including data beyond the $R_{500}$ radius, obtained from Planck observations of the Sunyaev–Zel’dovich (SZ) effect (adePlanckIntermediateResults2016).
@@
Comaa Density profile: $n_{e}(r)=n_{0}\left[1+\left(\frac{r}{r_{c}}\right)^{2}\right]^{-\frac{3\beta}{2}}$ $n_{0}=3.42\times 10^{-3}\mathrm{\penalty 10000\ cm}^{-3},r_{\mathrm{c}}=290\ \mathrm{kpc},\beta=0.75$
@@
Virgoc Density profile: $n_{e}(r)=n_{0,1}\left[1+\left(\frac{r}{r_{c,1}}\right)^{2}\right]^{-\frac{3\beta_{1}}{2}}\\ \quad\quad\quad+n_{0,2}\left[1+\left(\frac{r}{r_{c,2}}\right)^{2}\right]^{-\frac{3\beta_{2}}{2}}(r\lesssim 30\ kpc)\\ n_{\mathrm{e}}(r)=\frac{8.5\times 10^{-5}}{(r/\mathrm{Mpc})^{1.2}}\mathrm{\penalty 10000\ cm}^{-3}(r\gtrsim 250\ kpc)$ $n_{0,1}=8.9\times 10^{-2}\ \mathrm{cm}^{-3},r_{c,1}=4\ \mathrm{kpc},\beta_{1}=1,n_{0,2}=1.9\times 10^{-3}\ \mathrm{cm}^{-3},r_{c,2}=19.4\ \mathrm{kpc},\beta_{2}=0.47$
@@
    For Coma, We adopt the $\beta$ model (cavaliereDistributionHotGas1978a) for the density profile and a generalized Navarro Frenk White (nagaiTestingXRayMeasurements2007) model for pressure radial profile, with parameters taken from brielObservationComaCluster1992 and planckcollaboration:PlanckIntermediateResults2013a.
@@
    The functions and parameters used for the Perseus cluster in this study are consistent with those employed in consortiumProspectsGammaRay2023.
@@
    adePlanckIntermediateResults2016 combines a double-beta function for the inner region $(r\lesssim 30\ kpc)$ with a power-law function for the outer region $(r\gtrsim 250\ kpc)$ to describe the thermal electron number density distribution in the Virgo cluster. A simple linear interpolation was applied to the intermediate region. We converted the angular units into distance units.